Ticket Number: HEA-NOC/20080127-5 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver: HEA-NOC
Ticket Opened: 20080127 16:02 UTC Problem Start: 20080127 14:36 UTC
Ticket Update: 20080131 10:17 UTC Problem End: 20080131 08:15 UTC
Site/line: St. Angela's
There was a brief flap to St. Angela's (~4 seconds, only observed on
Traffic to St. Angela's.
20080127 15:22 CS Reported to BT, ref: 1-551828715
20080127 17:53 CS BT report seeing no alarms. Will monitor for another
20080128 15:38 LT No further issues for 25 hours. Closing ticket.
20080131 09:53 LT This link flapped several times again this morning,
between 08:09 and 08:15. There were 4 flaps lasting 10 seconds or less
(link did not report as down during these), and one link outage lasting
2.5 minutes. Contacting BT.
20080131 10:16 LT BT reference: 1-553173615.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20080127-5
HEAnet Limited. Registered in Ireland, No. 275301.